# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v2

@@
Figure 1: Comparison of LatentSeek with RL-based fine-tuning and Prompt Engineering. RL-based fine-tuning methods generally require iterative updates to model parameters guided by reward signals. Prompt engineering approaches depend heavily on manually designed prompts. In contrast, LatentSeek performs optimization within the latent space. Of note, the output of LatentSeek may be incoherent and semantically ungrounded ($\S$3.6).
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across a wide array of tasks, particularly in complex reasoning and deductive analysis (Brown et al., 2020, Chowdhery et al., 2022, OpenAI, 2023, Zhao et al., 2025). Despite these advancements, LLMs still exhibit difficulties in reasoning, particularly in tasks that demand structured thinking and meticulous step-by-step analysis (Wei et al., 2022, Kojima et al., 2022). Common approaches to enhancing the reasoning capabilities of LLMs involve training model parameters on reasoning-centric datasets or providing reasoning-oriented feedback (Ouyang et al., 2022, Bai et al., 2022, DeepSeek-AI, 2025). However, training methodologies such as supervised fine-tuning, reinforcement learning, and test-time training Sun et al. (2020), Hardt and Sun (2023) need to update model parameters, which incur substantial computational costs and present potential risks, including catastrophic forgetting of general competencies (Luo et al., 2025). Moreover, the widely used reinforcement learning approach may reduce the model’s exploration capacity Yue et al. (2025) and, in some cases, lead to the generation of overly verbose responses Aggarwal and Welleck (2025), Wu et al. (2025). In light of these challenges, we focus on an alternative paradigm, Test-Time Instance-Level Adaptation (TTIA), that does not require parameter updating and operates on a per-instance basis during the testing phase.
@@
Our innovative latent space TTIA method is simple yet surprisingly effective: Notably, LatentSeek achieves an average improvement of 10.75% over the CoT method on the GSM8K dataset, 3.93% on MATH-500, and 4.73% on AIME2024. Furthermore, when using LlaMA3.1-8B-Instruct as the backbone, LatentSeek surpasses prior arts including SimpleRL-Zoo Williams (1992) (+18.1%) and Genius Xu et al. (2025a) (+12.7%), on the GSM8K and MATH-500 datasets. To further assess the potential of the latent space, we conduct idealized experiments utilizing a perfect verifier that relies on ground truth and provides sparse rewards. The experimental results demonstrate a significant performance improvement, with an average gain of 19.12% points over CoT reasoning across all evaluated settings. This highlights the effectiveness of the knowledge elicitation mechanism employed by LatentSeek. Further experiments (see $\S$3.4) reveal that the model’s performance at test time improves proportionally with the number of update iterations. Notably, when equipped with an ideal verifier, LatentSeek is capable of elevating the performance of a 1.5B parameter model from 54.8% to 82.8% on the MATH-500 dataset, approaching the performance of OpenAI’s o1-preview model. These findings suggest that test-time scaling can be effectively implemented in the latent space, offering a viable alternative to conventional token-space scaling strategiesLiu et al. (2025), Yeo et al. (2025), Xu et al. (2025b).
@@ 2.2 Reasoning via Policy Gradient in Latent Space
where the expectation is approximated using the empirical mean in practical implementation. Please refer to Appendices˜C and D for a theoretical justification of this assumption and a detailed derivation of the policy-gradient update.
@@ 2.3 LatentSeek Algorithm
while $k\leq K$ and $r\leq\tau$ do $\triangleright$ $K$ is typically 1010
@@ 3 Empirical Results
In this section, we present our experimental setup, results, and analysis. Additional implementation details are provided in Sections˜F.1, F.3, F.2, F.4, F.5 and F.6.
@@ 3.1 Experimental Setup
In this subsection, we list our experimental setup.
Reward Mechanism.
We employ the mathematical reasoning prompts introduced by Lifshitz et al. (2025) to elicit the model’s self-reward computation. The same prompt structure is applied consistently across all tasks and models. For specific tasks and models, we further incorporate a format-based reward DeepSeek-AI (2025). To further illustrate the potential of our paradigm, we introduce a Perfect Sparse Reward Model (PSRM), the details of which are discussed in Section˜3.3.
Prompt Designation.
@@
Backbones.
In order to evaluate the generalizability of LatentSeek, we conduct experiments with pre-trained LLMs of different families and sizes: Qwen2-7B-Instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct Yang et al. (2024), Team (2024a), LLaMA3.1-8B-Instruct Team (2024b), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023).
Benchmarks.
Following Deng et al. (2024) and Liu et al. (2025), we focus on mathematical reasoning for evaluation. We conduct experiments on three datasets: GSM8K Cobbe et al. (2021), MATH-500 Hendrycks et al. (2021), and AIME2024. GSM8K and MATH-500 are commonly used for evaluating the performance of reasoning systems, thus providing a solid basis for comparative analysis. AIME2024, on the other hand, represents a more challenging dataset designed to assess the robustness and effectiveness of our proposed method in complex problem-solving scenarios.
Baselines.
@@
  • Prompting (Training-Free): CoT Wei et al. (2022) and Few-Shot CoT Lambert et al. (2025).
  • Explicit Search (Training-Free): Best-of-N (BoN) represents a highly effective search strategy, as demonstrated by Liu et al. (2025).
  • Reinforcement Learning: (1) Self Reward: Self-Rewarding Yuan et al. (2025), ScPO Prasad et al. (2024), CoH Liu et al. (2023a), and Genius Xu et al. (2025a). (2) Verifiable Reward: SimpleRL-Zoo Zeng et al. (2025), GRPO (GSM8K train set), and SPIN Chen et al. (2024a).
  • Latent Chain-of-Thought: Although iCoT Deng et al. (2024) requires an augmented training dataset for GSM8K Cobbe et al. (2021), it remains a computationally efficient baseline and is a canonical example of the latent CoT.
  • Supervised Fine-Tuning (SFT): Following Xu et al. (2025a), we apply SFT on Magpie 25K. For GSM8K, which includes a training set, we also report SFT performance on that set.
Table 1: Accuracy results (%) on GSM8K Cobbe et al. (2021), MATH-500 Hendrycks et al. (2021), and AIME2024 datasets. The best performances, excluding Perfect Sparse Reward, are highlighted in bold, and the second-best performances are underlined. Numbers highlighted in red indicate the performance gap relative to the CoT method. The symbol denotes methods that require parameter updates.
Methods Model Qwen2 Qwen2.5 LLaMA3.1 Mistral Avg
7B 1.5B 7B 14B 8B 7B
GSM8K: Prompt 1
CoT 68.01 68.08 88.48 92.03 50.19 23.28 65.01
BoN ($N=3$) 79.76 68.31 89.08 92.27 72.93 24.11 71.30
SFT 65.86 49.20 72.55 82.39 40.33 24.92 55.88
SFT (Magpie 25K) 76.50 66.48 83.01 90.30 70.81 13.72 66.80
iCoT Deng et al. (2024) 47.54 23.28 41.02 - 47.08 49.13 41.61
LatentSeek (Self) 84.38+16.37 69.37+1.29 89.46+0.98 92.49+0.46 78.54+28.35 23.88+0.60 73.02+8.01
LatentSeek (PSRM) 92.80+24.79 85.44+17.36 93.93+5.45 95.91+3.88 88.55+38.36 65.96+42.68 87.10+22.09
GSM8K: Prompt 2
CoT 65.20 15.31 66.41 91.81 69.07 41.70 58.25
BoN ($N=3$) 61.33 8.49 74.04 92.27 75.97 43.06 59.19
LatentSeek (Self) 80.21+15.01 44.20+28.89 85.06+18.65 92.72+0.91 83.70+14.63 44.58+2.88 71.74+13.49
LatentSeek (PSRM) 92.80+27.60 67.48+52.17 93.78+27.37 96.13+4.32 94.77+25.70 78.24+36.54 87.20+28.95
MATH-500: Prompt 1
CoT 51.40 54.80 72.80 77.20 47.60 16.40 53.37
BoN ($N=3$) 53.40 47.40 75.40 78.80 51.20 16.60 53.80
SFT (Magpie 25K) 46.6 44.40 55.40 68.2 31.00 3.2 41.46
LatentSeek (Self) 57.40+6.00 55.60+1.80 75.60+2.80 80.00+2.80 54.60+7.00 17.20+0.60 56.73+3.36
LatentSeek (PSRM) 75.80+24.40 75.80+21.00 86.40+13.60 87.20+10.00 74.60+27.00 41.80+25.40 73.60+20.23
MATH-500: Prompt 2
CoT 37.40 27.60 53.80 68.00 40.40 8.20 39.23
BoN ($N=3$) 41.60 24.80 55.80 64.20 44.40 11.60 40.40
LatentSeek (Self) 44.80+7.39 32.20+4.60 57.60+3.80 71.00+3.00 47.00+6.60 9.80+1.60 43.73+4.50
LatentSeek (PSRM) 67.80+30.40 57.60+30.00 77.20+23.40 86.00+18.00 66.40+26.00 31.40+23.20 64.39+25.16
AIME2024: Prompt1
CoT 0.0 3.33 6.67 10.00 0.00 0.00 3.33
BoN ($N=3$) 0.00 0.00 10.00 16.67 0.00 0.00 4.45
SFT (Magpie 25K) 3.33 0.0 3.33 10.00 3.33 0.00 3.33
LatentSeek (Self) 3.33+3.33 6.67+3.33 13.33+6.37 16.67+6.67 10.00+10.00 3.33+3.33 8.33+4.45
LatentSeek (PSRM) 13.33+13.33 6.67+3.33 16.67+10.00 26.67+16.67 10.00+10.00 3.33+3.33 12.78+9.45
AIME2024: Prompt2
CoT 0.00 0.00 0.00 3.33 0.00 0.00 0.56
BoN ($N=3$) 3.33 0.00 6.67 10.00 6.67 0.00 4.45
LatentSeek (Self) 3.33+3.3 3.33+3.3 13.33+13.33 10.00+6.67 6.67+6.67 0.00+0.00 5.55+5.00
LatentSeek (PSRM) 3.33+3.33 6.67+6.67 13.33+13.33 23.33+20.00 10.00+10.00 0.00+0.00 9.44+8.88
Table 2: Accuracy score (%) compared with more baseline methods on GSM8K and MATH-500 datasets with Llama3.1-8B as backbone.
@@
Few-Shot CoT Lambert et al. (2025) - Instruct 83.4 42.5 63.0
Self-Rewarding Yuan et al. (2025) Self Instruct 76.0 30.2 53.1
ScPO Prasad et al. (2024) Self Instruct 71.1 31.00 51.1
SPIN Chen et al. (2024a) Data Instruct 74.9 31.5 53.2
CoH Liu et al. (2023a) Self Instruct 74.4 32.3 53.3
Genius Xu et al. (2025a) Self Instruct 78.3 34.6 56.5
iCoT Deng et al. (2024) Data Instruct 47.1 - 47.1
BoN ($N=3$) Self Instruct 76.0 51.2 63.6
@@
SimpleRL-Zoo Zeng et al. (2025) Data Base 79.2 23.0 51.1
@@ 3.2 State-of-the-art Test-time Reasoning Performance
Best Performance on GSM8K Cobbe et al. (2021) and MATH-500 Hendrycks et al. (2021).
As demonstrated in Table˜2, using LLaMA3.1-8B-Instruct as the backbone, our method outperforms all baseline approaches across both the GSM8K and MATH-500 datasets. Specifically, on the GSM8K dataset, our method achieves an improvement of 14.6% points over the CoT baseline, and a 7.0%-point gain on the MATH-500 dataset. In comparison to the training-free BoN approach, our method yields improvements of 7.7% points and 3.4% points on GSM8K and MATH-500, respectively. When compared to the SFT (Magpie 25K) Xu et al. (2025a), the minimum improvement across both datasets is 12.9% points. Moreover, our approach surpasses the state-of-the-art self-reward RL-based method, Genius Xu et al. (2025a), by 5.4% points on GSM8K and 20.0 points on MATH-500. Notably, it also outperforms iCoT Deng et al. (2024)—a latent method trained on an augmented GSM8K dataset—by 36.6% points on the GSM8K test set. Even when compared to reinforcement learning with verifiable feedback methods, like SimpleRL-Zoo Zeng et al. (2025), which was trained on the base version of the model, our method exceeds it by an average of 9.1% points across both datasets. These results underscore the effectiveness of our approach in tackling reasoning tasks.
@@
AIME2024 results are listed in Table˜1. AIME2024 is a highly challenging mathematics benchmark consisting of only 30 problems, designed to evaluate advanced reasoning capabilities. Experimental results demonstrate the power and potential of our method in complex settings. As shown in Table˜1, our approach consistently outperforms all baselines, achieving an average improvement of 4.73% points over CoT across all model families and prompt configurations. Notably, even weaker models such as Mistral and Qwen2.5-1.5B-Instruct benefit from our method, with average gains of 1.67% and 3.33% points compared with CoT, respectively. The largest improvement is observed on AIME2024 (prompt 1), where our method surpasses the CoT baseline by 13.33% points and outperforms BoN by 6.67% points using Qwen2.5-7B-Instruct.
Generalizability Across Backbones.
As presented in Section˜3, the generalizability of our approach with respect to backbone models can be examined along two dimensions: (1) Model Families: LatentSeek demonstrates superior performance across multiple model families. Specifically, it achieves the best results when using backbones from the Qwen2, Qwen2.5, and LLaMA3.1 series. Additionally, it ranks first in four out of six evaluation settings and second in one when employing a relatively weaker backbone from the Mistral family. The suboptimal performance of the Mistral-based model is attributed to its limited capacity to function effectively as a reward model. (2) Model Size: At the 1.5B parameter scale, our method consistently outperforms all baseline models across diverse datasets and prompt types, with a particularly notable gain of 28.89% points over the CoT baseline on the GSM8K dataset (Prompt 2). At the 7–8B scale, models based on Qwen2, Qwen2.5, and LLaMA3.1 consistently surpass all baselines. In particular, Qwen2 yields an average improvement of 11.75% points over BoN on GSM8K. At the 14B scale, our approach continues to outperform all baselines, achieving a significant 6.67%-point improvement on the AIME2024 dataset. These results provide robust evidence of our method’s generalizability across diverse model families and scales.
Generalizability on Model-specific prompt.
The Qwen2.5 series was explicitly trained using Prompt 1 Team (2024a); nevertheless, our methods still achieve notable performance gains. Specifically, the proposed method attains average scores of 83.77% on GSM8K and 70.40% on MATH-500, reflecting improvements of 0.91% and 2.47% points, respectively, despite the strong baseline performance achieved by CoT. Additionally, significant improvements are consistently observed across all other models utilizing Prompt 1, apart from Qwen2.5. For example, Qwen2 achieves substantial increases of 16.37% points on GSM8K and 7.39% points on MATH-500 compared to the CoT baseline. These results demonstrate that our method effectively enhances performance not only with normal prompts but also with model-specific ones.

### v3

@@
Figure 1: Comparison of LatentSeek with RL-based fine-tuning and Prompt Engineering. RL-based fine-tuning methods generally require iterative updates to model parameters guided by reward signals. Prompt engineering approaches depend heavily on manually designed prompts. In contrast, LatentSeek performs optimization within the latent space. Of note, the output of LatentSeek may be incoherent and semantically ungrounded ($\S$3.8).
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated exceptional performance across a wide array of tasks, particularly in complex reasoning and deductive analysis (Brown et al., 2020, Chowdhery et al., 2023, OpenAI, 2023, Zhao et al., 2025). Despite these advancements, LLMs still exhibit difficulties in reasoning, particularly in tasks that demand structured thinking and meticulous step-by-step analysis (Wei et al., 2022, Kojima et al., 2022). Common approaches to enhancing the reasoning capabilities of LLMs involve training model parameters on reasoning-centric datasets or providing reasoning-oriented feedback (Ouyang et al., 2022, Bai et al., 2022, DeepSeek-AI, 2025). However, training methodologies such as supervised fine-tuning, reinforcement learning, and test-time training Sun et al. (2020), Hardt and Sun (2024) need to update model parameters, which incur substantial computational costs and present potential risks, including catastrophic forgetting of general competencies (Luo et al., 2025). Moreover, the widely used reinforcement learning approach may reduce the model’s exploration capacity Yue et al. (2025) and, in some cases, lead to the generation of overly verbose responses Aggarwal and Welleck (2025), Wu et al. (2025). In light of these challenges, we focus on an alternative paradigm, Test-Time Instance-Level Adaptation (TTIA), that does not require parameter updating and operates on a per-instance basis during the testing phase.
@@
Our innovative latent space reasoning method is simple yet effective: Notably, LatentSeek yields average gains of 15.23 points over BoN on GSM8K, 4.72 points over Self-Reflection on MATH-500, and 6.67 points over CoT on AIME2024 (Table˜1). Furthermore, when using LLaMA3.1-8B-Instruct as the backbone, LatentSeek surpasses prior arts including SimpleRL-Zoo (Williams, 1992) (+18.1) and Genius (Xu et al., 2025a) (+12.7), on the GSM8K and MATH-500 datasets (Table˜2). Further experiments ( $\S$3.4) show that test-time performance improves with the number of update iterations, suggesting a complementary scaling axis: the number of optimization steps in latent space. To better characterize the latent space, we conduct idealized experiments using a perfect, ground-truth verifier that issues sharp rewards. Across all evaluation settings, this yields an average gain of 19.12 points over CoT (Table˜3), highlighting the benefits of pure exploration in the latent space. Apart from the above, our case studies ($\S$3.8) reveal that LatentSeek often attains the correct answer even when its intermediate reasoning is unintelligible to human readers. This divergence reveals that the optimal reasoning path for LLMs need not mirror human reasoning, and that optimization in latent space may be a more native and effective paradigm.
@@ 2 Test-Time Instance-Level Policy Gradient in Latent Space
In this section, we begin with a test-time, instance-level reasoning formulation and then derive our own formulation, which subsequently motivates the policy-gradient method in latent space.
@@ 2.2 Reasoning via Policy Gradient in Latent Space
where the expectation is approximated using the empirical mean in practical implementation. The independence assumption is adopted for two main reasons. (1) Enlarged Exploration Space. Without this assumption, the autoregressive structure would force the optimization to concentrate on the first latent token, since all subsequent tokens are conditioned on it. This collapses the effective search space. By contrast, imposing independence decouples the latent components, yielding a substantially larger exploration space and a more flexible “launch pad” for generating the subsequent reasoning trajectory $\prod_{t=N+1}^{T}\pi(x_{t}\mid x_{<t},\mathbf{c})$. (2) Theoretical Guidance. Our analysis in Appendix˜C draws an analogy to Multi-Prover Interactive Proofs (MIP), and shows that the potential loss introduced by the independence assumption can be mitigated by both the base model’s ability to faithfully generate the follow-up reasoning trajectory and the reward model’s ability to evaluate it. Please refer to Appendices˜C and D for a theoretical guidance of this assumption and a detailed derivation of the policy-gradient update.
@@ 2.3 LatentSeek Algorithm
while $k\leq K$ and $r\leq\tau$ do $\triangleright$ $K$ is typically $10$
@@ 3 Empirical Results
In this section, we perform empirical results to validate our LatentSeek algorithm.
Table 1: Accuracy results. Bold: best performance, Underlined: second-best. Numbers highlighted in red indicate the performance gap relative to the CoT method. Each rollout baseline uses a rollout number comparable to LatentSeek ’s update iterations for fairness. See Section˜E.1 for larger-number comparisons, Section˜E.3 for token consumption, Section˜E.8 for performance trends across different fraction ratios, and Figure˜4 for results of Qwen3 on AIME2024. : requires training.
Methods Model Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8K: Prompt 1
CoT 68.01 68.08 88.86 92.03 50.19 73.43
BoN 79.76 68.31 89.08 92.27 72.93 80.47
Self-Reflection (Shinn et al., 2023) 67.85 68.84 88.48 92.21 52.16 73.91
Self-Consistency 73.24 69.07 89.23 92.19 55.34 75.81
SFT 65.86 49.20 72.55 82.39 40.33 62.07
SFT (Magpie 25K) 76.50 66.48 83.01 90.30 70.81 77.42
iCoT (Deng et al., 2024) 47.54 23.28 41.02 - 47.08 39.73
LatentSeek (Self) 84.38+16.37 70.89+2.81 90.14+1.28 92.49+0.46 78.54+28.35 83.29+9.86
GSM8K: Prompt 2
CoT 65.20 15.31 66.41 91.81 69.07 61.56
BoN 61.33 6.14 74.04 92.27 75.97 61.95
Self-Reflection (Shinn et al., 2023) 68.16 27.98 72.63 91.96 76.22 67.39
Self-Consistency 64.37 8.79 69.22 92.49 74.30 61.83
LatentSeek (Self) 80.21+15.01 44.20+28.89 85.06+18.65 92.72+0.91 83.70+14.63 77.18+15.62
MATH-500: Prompt 1
CoT 51.40 54.80 72.80 77.20 47.60 60.76
BoN 53.40 47.40 75.40 78.80 51.20 61.24
Self-Reflection (Shinn et al., 2023) 53.00 53.00 72.40 75.80 45.40 59.92
Self-Consistency 53.60 51.60 73.20 77.40 50.40 61.24
SFT (Magpie 25K) 46.60 44.40 55.40 68.20 31.00 49.12
LatentSeek (Self) 57.40+6.00 55.60+0.80 75.60+2.80 80.00+2.80 54.60+7.00 64.64+3.88
MATH-500: Prompt 2
CoT 37.40 29.40 53.80 68.00 40.40 45.80
BoN 41.60 29.40 55.80 64.20 44.40 47.08
Self-Reflection (Shinn et al., 2023) 43.00 32.20 55.60 70.00 35.20 47.20
Self-Consistency 43.20 17.20 57.00 65.80 45.20 45.68
LatentSeek (Self) 44.80+7.40 34.40+5.00 57.60+3.80 71.00+3.00 47.00+6.60 50.96+5.16
AIME2024: Prompt1
CoT 0.00 3.33 6.67 10.00 0.00 4.00
BoN 0.00 0.00 10.00 16.67 0.00 5.33
Self-Reflection (Shinn et al., 2023) 0.00 6.67 6.67 10.00 0.00 4.67
Self-Consistency 3.33 0.00 3.33 13.33 6.67 5.33
SFT (Magpie 25K) 3.33 0.00 3.33 10.00 3.33 4.00
LatentSeek (Self) 3.33+3.33 6.67+3.33 13.33+6.67 16.67+6.67 10.00+10.00 10.00+6.00
AIME2024: Prompt2
CoT 0.00 0.00 0.00 3.33 0.00 0.67
BoN 3.33 0.00 6.67 10.00 6.67 5.33
Self-Reflection (Shinn et al., 2023) 0.00 3.33 0.00 6.67 6.67 3.33
Self-Consistency 3.33 0.00 3.33 3.33 0.00 2.00
LatentSeek (Self) 3.33+3.33 3.33+3.33 13.33+13.33 10.00+6.67 6.67+6.67 7.33+6.67
@@ 3.1 Experimental Setup
In this subsection, we list our experimental setup, including reward mechanism, prompt designation, backbones, benchmarks, and baselines. See Sections˜F.1, F.3, F.2, F.4, F.5 and F.6 for more details.
Reward Mechanism
We employ the mathematical reasoning prompts introduced by Lifshitz et al. (2025) to elicit the model’s self-reward computation. The reward prompts remain consistent across all datasets. For some models on specific tasks, we further incorporate a format-based reward (DeepSeek-AI, 2025) to enhance their instruction-following capability. To further analysis the potential of our paradigm, we introduce a Perfect Sparse Reward Model (PSRM), the details of which are discussed in $\S$3. All generated reward is chosen to be a number between -1 and 0.
Prompt Designation
@@
Backbones
To evaluate the generalizability of our approach, we conduct experiments with pre-trained LLMs of different families and sizes: Qwen2-7B-Instruct, Qwen2.5-1.5B, 7B, 14B-Instruct (Yang et al., 2024a, Team, 2025), LLaMA3.1-8B-Instruct (Team, 2024).
Benchmarks
Following Deng et al. (2024), Liu et al. (2025), we focus on mathematical reasoning for evaluation. We evaluate on two standard benchmarks for reasoning, GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021), and a harder dataset AIME2024.
Baselines
@@
  • Prompting (Training-Free): CoT (Wei et al., 2022) and Few-Shot CoT (Lambert et al., 2024).
  • Explicit Search (Training-Free): Best-of-N (BoN) represents a highly effective search strategy, as demonstrated by Liu et al. (2025). Self-Reflection (Shinn et al., 2023) is a search strategy imitating humans’ reasoning pattern. Self-Consistency, also named majority vote, is a search mechanism in explicit space that stems from a multi-agent system.
  • Reinforcement Learning: (1) Self Reward: Self-Rewarding (Yuan et al., 2024), ScPO (Prasad et al., 2025), CoH (Liu et al., 2024b), and Genius (Xu et al., 2025a). (2) Verifiable Reward: SimpleRL-Zoo (Zeng et al., 2025), GRPO (GSM8K train set), and SPIN (Chen et al., 2024b).
  • Latent Chain-of-Thought: Although iCoT (Deng et al., 2024) requires an augmented training dataset for GSM8K (Cobbe et al., 2021), it remains a canonical example of the latent CoT.
  • Supervised Fine-Tuning (SFT): Following (Xu et al., 2025a), we apply SFT on Magpie 25K. For GSM8K, we also report SFT performance using its provided training set.
Sampling Method
We use greedy decoding to sample from LatentSeek, as it is more computationally efficient and more robust. (See Section˜E.4 for an experimental comparison).
Table 2: Accuracy score compared with more baseline methods on GSM8K and MATH-500 datasets with Llama3.1-8B as backbone. The best performances are highlighted in bold.
@@
Few-Shot CoT (Lambert et al., 2024) - Instruct 83.4 42.5 63.0
BoN Self Instruct 76.0 51.2 63.6
Self-Rewarding (Yuan et al., 2024) Self Instruct 76.0 30.2 53.1
Self-Relfection (Shinn et al., 2023) Self Instruct 76.2 45.4 60.8
Self-Consistency Self Instruct 55.3 48.2 51.8
ScPO (Prasad et al., 2025) Self Instruct 71.1 31.0 51.1
CoH (Liu et al., 2024b) Self Instruct 74.4 32.3 53.3
Genius (Xu et al., 2025a) Self Instruct 78.3 34.6 56.5
SPIN (Chen et al., 2024b) Data Instruct 74.9 31.5 53.2
iCoT (Deng et al., 2024) Data Instruct 47.1 - 47.1
@@
SimpleRL-Zoo (Zeng et al., 2025) Data Base 79.2 23.0 51.1
@@ 3.2 State-of-the-art Test-time Reasoning Performance
Best Performance on GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021).
As shown in Table˜1, LatentSeek surpasses every baseline across all 30 settings and achieves the best average performance over all datasets. When averaged across backbones, the improvement over BoN is 15.23 points on GSM8K and against Self-Reflection is 4.72 points on MATH-500. Compared with Self-Consistency, the gap is 15.35 on GSM8K and 5.28 on MATH-500. The Qwen2.5 series was explicitly trained with Prompt-1–distilled data (Team, 2025), yielding strong CoT performance; nevertheless, LatentSeek achieves further improvements and attains the highest scores across baselines. Using LLaMA3.1 as the backbone, as shown in Table˜2, we compare against a broader set of baselines. On GSM8K, our method surpasses the CoT baseline by 14.6 points and BoN by 7.7 points; on MATH-500, the improvements are 7.0 and 3.4 points, respectively. It also outperforms the self-reward RL method Genius (Xu et al., 2025a) by 5.4 points on GSM8K and 20.0 points on MATH-500. Even compared with verifiable-feedback RL methods such as SimpleRL-Zoo (Zeng et al., 2025), trained on the base model, LatentSeek leads by an average of 18.1 points across the two datasets.
Robustness Across Backbones.
The robustness of our method across backbones can be evaluated along two axes: (1) Model Families: LatentSeek delivers best results across different models. Focusing on 7–8B models and using the same averaging protocol over GSM8K, MATH-500, and both prompts, LatentSeek surpasses BoN by 7.67 points on Qwen2, exceeds Self-Reflection by 4.82 on Qwen2.5, and improves over Self-Consistency by 9.65 on LLaMA3.1. (2) Model Size: At 1.5B parameters, LatentSeek surpasses BoN on GSM8K by 28.89 points and at 14B parameters, the margin on MATH-500 attains 6.80 points. These results provide robust evidence of our method’s robustness across diverse model families and scale.
@@
AIME2024 results appear in Table˜1. Our method consistently outperforms all baselines, achieving an average gain of 6.34 points over CoT across model families and prompt configurations. The largest improvement occurs on AIME2024 with Prompt 2, where our approach exceeds CoT by 13.33 points and outperforms BoN by 6.67 points using Qwen2.5-7B-Instruct. Averaged over backbones, LatentSeek also surpasses BoN by 4.67 points, Self-Reflection by 5.33 points, and Self-Consistency by 4.67 points, indicating strong effectiveness in more complex settings. For the performance of a stronger model (Qwen3-4B-Instruct-2507 (Yang et al., 2025)) on AIME2024, please refer to Section˜E.2.